OpenStudy (anonymous):

The variables x and y vary directly. Write an equation that relates x and y, when x = 4 and y = 2.

OpenStudy (anonymous):

what do you mean by vary directly?

OpenStudy (anonymous):

not sure exactly, it's part of the problem.

OpenStudy (agreene):

y=x-2 plug in to check: (4,2) 2=4-2 2=2 atleast, I think that is what it is asking for O.o

OpenStudy (anonymous):

A. y = -2x B. y = -1/2x C. y = 1/2x D. y = 2x

OpenStudy (anonymous):

so A?